OpenStudy (anonymous):

the length of a rectangle, I, is six meters more than its width, w. if the perimeter of the rectangle is 36 meters, what is the length of the rectangle, in meters?

OpenStudy (anonymous):

Using P as perimeter, l as length and W and width: For a rectangle: \[P = 2 l + 2 w\] But they tell us that the lenght is six meters more than its width. We write that in this manner: \[l = w+6\] If we replace the first l for this one, we will end up with \[P = 2 (w+6) + 2w\] \[P = 2 w+12 + 2w\] \[P = 4w + 12\] And since we know the perimeter is 36, \[36 = 4w + 12\] \[4w = 24\] \[w = 6\] But they ask us the length. So going back, we remember that l = w + 6. Since w is 6, l must be 12.

OpenStudy (anonymous):

so its A